Two printing presses, working together, can complete a job in 2 hours. If one press requires 6 hours to do the job alone, how many hours would the second press ned to complete the job alone?
Answer by ptaylor(2048)

Let x=amount of time required for the second press to complete the job alone
Then the second press works at the rate of 1/x of the job per hour
Together, the two presses work at the rate of 1/2 of the job per hour
And the first press works at the rate of 1/6 of the job per hour
So, our equation to solve is:
1/6+1/x=1/2 multiply each term by 6x
x+6=3x subtract x from each side
x-x+6=3x-x collect like terms
2x=6 divide both sides by 2
x=3 number of hours needed for the second press to complete the job
CK
1/6+1/3=1/2
1/6+2/6=1/2
3/6=1/2
1/2=1/2
also
(1/6)*2=1/3
(1/3)*2=2/3
2/3+1/3=1 (1 job, that is)
1=1
Hope this helps---ptaylor